OpenStudy (anonymous):

i already have an answer. i just need a comparison. What is the solution set for |3x – 24| 27? what u guys get?

OpenStudy (anonymous):

What did you get?

OpenStudy (saifoo.khan):

|3x – 24| = 27

OpenStudy (saifoo.khan):

is this is it right?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

3x - 24 = 27 & 3x - 24 = -27

OpenStudy (anonymous):

or actually, it cannot be both.

OpenStudy (saifoo.khan):

Oh, i mean OR.

OpenStudy (saifoo.khan):

3x - 24 = 27 OR 3x - 24 = -27

OpenStudy (anonymous):

wait nvrmnd. its rather greater than or equal to

OpenStudy (anonymous):

27

OpenStudy (saifoo.khan):

:O what?

OpenStudy (anonymous):

blamb, did you get a solution?

OpenStudy (anonymous):

its not an equal sign. its rather greater than or equals to

OpenStudy (anonymous):

yes i did

OpenStudy (saifoo.khan):

try inserting the values, if it gets the solutions, then your answers are crrct.

OpenStudy (anonymous):

indeed. If you have solutions you can plug it in and see if the expression is true. if not, you've got the wrong solutions.